# Vehicle Top Tag Assisted Vehicle-Road Cooperative Localization For Autonomous Public Buses

### v1

@@ I Introduction
Accurate (i.e. centimetre-level) vehicle localization or for short vehicle localization is indispensable to autonomous vehicles. State-of-the-art (SOTA) vehicle localization systems normally rely on certain exteroceptive sensors such as GNSS, LiDAR, and vision system (or camera), augmented by proprioceptive sensors such as IMU. Relevant methods can be mainly categorized into GNSS based ones, LiDAR based ones, and vision based ones. These categories of vehicle localization methods are not mutually exclusive. In fact, merits of multi-sensor fusion architectures have long since been advocated [1] [2] [3]. One may refer to [4] for a knowledge of vehicle localization methods in early years and refer to [5] for a survey of relevant methods in recent years.
@@
Research fruits flourish in literature, yet no existing method can satisfy all the feasibility, reliability, explainability, society and economy concerns for our dedicated “vehicle-road-cloud” projects, the core part of which consists in running autonomous public buses  [^1] Qingfei.AI is dedicated to research, development, and manufacturing of social intelligent systems such as autonomous public buses. in complicated application scenarios such as illustrated in Fig. 1.
@@
To handle difficult intersection scenarios, a technically sound idea is to add special road marks [13] [14] at intersections for guidance. But such kind of methods have difficulty in satisfying society concerns. First, road marks are strictly controlled by the government’s road administration bureau. Anyone and even the bureau itself have no right to arbitrarily add any road mark especially alien road mark  [^3] Special road marks or tags usually look alien or strange to humans.. Second, alien road marks may cause confusion to human drivers and cause traffic disorder. Third, alien road marks may give people a feeling of visual intrusion and are socio-psychologically undesirable. The last two reasons somehow account for the first reason namely why road marks in traffic environment should be strictly controlled by the government.
Another technically sound idea is to take advantage of the V2X strategy which has been finding its way in more and more “vehicle-road-cloud” projects in recent years. Multi-vehicle cooperative localization [15] [16] can be valuable to handle difficult intersection scenarios. However, according to our own experiences, difficulty of putting this into practice lies rather at policy and administration levels than at the technique level. Simply speaking, why should other vehicles cooperate with the autonomous public bus? After all, the government has no right to force social vehicles to do.
@@ IV Vehicle top tag pose estimation
For an AprilTag, its control points are its four outer corners, as illustrated by the circle, plus, star marks namely $\circ$”, “$+$”, “*” in Fig. 5. The distinct marks are shown to convey that the order of tag control points can also be known via AprilTag detection. Positions of the four physical corners of the AprilTag in the vehicle top tag coordinates system  [^6] A conventional way of establishing the vehicle coordinates system is to let the origin be located at the vehicle center projection on the ground plane and let the $x$-axis, $y$-axis, $z$-axis face forward, left-side, upward respectively. The vehicle top tag coordinates system differs from the conventional vehicle coordinates system only by a vehicle height, with its origin located at the vehicle center projection on the vehicle top tag plane instead of the ground plane. Since autonomous navigation depends on the vehicle horizontal pose and the two coordinates systems have no difference in their horizontal poses, for analysis convenience, we set the vehicle coordinates system identical to the vehicle top tag coordinates system. Whenever vehicle top tag pose estimation is mentioned throughout this paper, vehicle pose estimation (i.e. vehicle localization) is mentioned implicitly as well. can be known a priori. For example, given a single AprilTag installed at the vehicle top center (see Fig. 9-top-left) and the AprilTag width $2w$, then the four physical corners always have the fixed vehicle top tag coordinates
@@ IV-A Basic version of vehicle top tag pose estimation
where :=assign:=:= denotes value assignment rather than equality, $\mathbf{R}(\cdot)$ denotes the rotation function that maps a rotation vector to its associated rotation matrix, and $\mathbf{R}^{-1}(\cdot)$ denotes the inverse rotation function that maps a (quasi) rotation matrix to its associated rotation vector.
@@ V-A Performance with moderate (low) resolution
[…]
@@ VI Conclusion
In our application context, we focus on handling difficult intersection scenarios for autonomous public buses, yet the proposed methodology of vehicle top tag assisted localization can be applied in much wider way. First, on the side of vehicles, the proposed methodology can be directly extended to practical applications concerning automation of other kinds of big vehicles  [^10] Yet we have to admit that we had better refrain from extending the proposed methodology too generally to automation of small vehicles such as taxis, because special tags on their tops are in people’s normal views in traffic environment and may give people a feeling of visual intrusion that is socio-psychologically undesirable.. Second, on the side of roads, the proposed methodology can be directly extended to other road scenarios wherever visual RSUs are available, not limited only to intersection scenarios.

### v2

@@ I Introduction
Accurate (i.e. centimetre-level) vehicle localization or for short vehicle localization is indispensable to autonomous vehicles. State-of-the-art (SOTA) vehicle localization systems normally rely on certain exteroceptive sensors such as GNSS, LiDAR, and vision system (or camera), augmented by proprioceptive sensors such as IMU. Relevant methods can be mainly categorized into GNSS based ones, LiDAR based ones, and vision based ones. These categories of vehicle localization methods are not mutually exclusive. In fact, merits of multi-sensor fusion architectures have long since been advocated [1] [2] [3] [4]. One may refer to [5] for a knowledge of vehicle localization methods in early years and refer to [6] for a survey of relevant methods in recent years.
@@
Research fruits flourish in literature, yet no existing method can satisfy all the feasibility, reliability, explainability, society and economy concerns for our dedicated “vehicle-road-cloud” projects, the core part of which consists in running autonomous public buses in complicated application scenarios such as illustrated in Fig. 1.
@@
To handle difficult intersection scenarios, a technically sound idea is to add special road marks [14] [15] [16] at intersections for guidance. But such kind of methods have difficulty in satisfying society concerns. First, road marks are strictly controlled by the government’s road administration bureau. Anyone and even the bureau itself have no right to arbitrarily add any road mark especially alien road mark  [^2] Special road marks or tags usually look alien or strange to humans.. Second, alien road marks may cause confusion to human drivers and cause traffic disorder. Third, alien road marks may give people a feeling of visual intrusion and are socio-psychologically undesirable. The last two reasons somehow account for the first reason namely why road marks in traffic environment should be strictly controlled by the government.
Another technically sound idea is to take advantage of the V2X strategy which has been finding its way in more and more “vehicle-road-cloud” projects in recent years. Multi-vehicle cooperative localization [17] [18] [19] [20] can be valuable to handle difficult intersection scenarios. However, according to our own experiences, difficulty of putting this into practice lies rather at policy and administration levels than at the technique level. Simply speaking, why should other vehicles cooperate with the autonomous public bus? After all, the government has no right to force social vehicles to do so.
@@ IV Vehicle top tag pose estimation
For an AprilTag, its control points are its four outer corners, as illustrated by the circle, plus, star marks namely $\circ$”, “++”, “*” in Fig. 5. The distinct marks are shown to convey that the order of tag control points can also be known via AprilTag detection. Positions of the four physical corners of the AprilTag in the vehicle top tag coordinates system  [^5] A conventional way of establishing the vehicle coordinates system is to let the origin be located at the vehicle center projection on the ground plane and let the $x$-axis, $y$-axis, $z$-axis face forward, left-side, upward respectively. The vehicle top tag coordinates system differs from the conventional vehicle coordinates system only by a vehicle height, with its origin located at the vehicle center projection on the vehicle top tag plane instead of the ground plane. Since autonomous navigation depends on the vehicle horizontal pose and the two coordinates systems have no difference in their horizontal poses, for analysis convenience, we set the vehicle coordinates system identical to the vehicle top tag coordinates system. Whenever vehicle top tag pose estimation is mentioned throughout this paper, vehicle pose estimation (i.e. vehicle localization) is mentioned implicitly as well. can be known a priori. For example, given a single AprilTag installed at the vehicle top center (see Fig. 9-top-left) and the AprilTag width $2w$, then the four physical corners always have the fixed vehicle top tag coordinates
@@ IV-A Basic version of vehicle top tag pose estimation
where :=:= denotes value assignment rather than equality, $\mathbf{R}(\cdot)$ denotes the rotation function that maps a rotation vector to its associated rotation matrix, and $\mathbf{R}^{-1}(\cdot)$ denotes the inverse rotation function that maps a (quasi) rotation matrix to its associated rotation vector.
@@ V-A Performance with moderate (low) resolution
Performance statistics demonstrated in Fig. 6 are obtained under the assumption that the bus has no height disturbance. In practice, the autonomous public bus indeed has moderate height disturbance especially when passing intersections  [^7] Traffic rules require that vehicles move at low speeds ($\leq 30\mbox{ km/h}$) when passing intersections. Moving at low speeds, vehicles especially buses which normally have good suspension systems tend to have rather stable height.. On the other hand, to test robustness of H.Opt. and S.Opt. against height disturbance, we intentionally exaggerate the height disturbance to $10$ centimetres. Relevant performance statistics are demonstrated in Fig. 7.
@@ VI Conclusion
In our application context, we focus on handling difficult intersection scenarios for autonomous public buses, yet the proposed methodology of vehicle top tag assisted localization can be applied in much wider way. First, on the side of vehicles, the proposed methodology can be directly extended to practical applications concerning automation of other kinds of big vehicles. Second, on the side of roads, the proposed methodology can be directly extended to other road scenarios wherever visual RSUs are available, not limited only to intersection scenarios.